### v2

# Prime and Co-prime Integer Matrices

Abstract
This paper investigates prime and co-prime integer matrices and their properties. It characterizes all pairwise co-prime integer matrices that are also prime integer matrices. This provides a simple way to construct families of pairwise co-prime integer matrices, that may have applications in multidimensional co-prime sensing and multidimensional Chinese remainder theorem.
Index Terms:
Prime integer matrices, Gaussian integers, Gaussian primes, co-primality, Chinese remainder theorem (CRT), co-prime sensing.
I Introduction
It is well-known that prime integers are the most important subjects in mathematics and have played vital roles in applications, such as cryptography, digital communications, and signal processing etc. It plays the key role in the Chinese remainder theorem (CRT) that is to reconstruct a large integer from its remainders modulo several small moduli and has many applications in cryptography, digital communications and signal processing as well [1]. For a given set of moduli, the largest integer that can be uniquely reconstructed from its remainders is smaller than the least common multiple (lcm) of all the moduli and the lcm achieves the maximal if all the moduli are pairwise co-prime, when the moduli sizes are upper bounded. This means that it is critical to have a large size set of pairwise co-prime integers.
The conventional prime integers have been generalized to other prime algebraic integers in algebraic number fields with the corresponding CRT [3, 10, 11]. One example is Gaussian integers that are complex numbers with the conventional integer real and imaginary parts. Prime Gaussian integers are called Gaussian primes with simple and well-known necessary and sufficient conditions in terms of the conventional prime integers (called rational primes) [3]. In fact, an algebraic integer can be equivalently represented by an integer matrix where all the elements are the conventional integers. For a quadratic algebraic number field, see, for example, [10, 11], and in particular for a Gaussian integer of real and imaginary parts $a$ and $b$, respectively, its equivalent $2\times 2$ integer matrix representation is $\left(\begin{array}[]{cc}a&-b\\ b&a\end{array}\right)$. Thus, a Gaussian prime also means that its equivalent $2\times 2$ integer matrix (representation) is prime. Similarly, the CRT for Gaussian integers can be thought of as a CRT for $2\times 2$ integer matrices of the above special form, where Gaussian primes play the key role as well. It has recent applications in multi-channel self-reset analog-to-digital converter (SR-ADC) using modulo sampling for a complex valued bandlimited signal [18, 19].
Prime integers have also been generalized to prime integer matrices in [2] where it obtains that an integer matrix is prime if and only if its determinant absolute value is a conventional prime integer. Recently, CRT has been generalized to general integer matrices of any form and any dimension in [15, 16] to reconstruct integer vectors from their vector remainders modulo several integer matrices called matrix moduli, i.e., multidimensional CRT (MD-CRT). Similar to the conventional CRT for integers, to have a large range of uniquely reconstructable integer vectors from their vector remainders, it is critical to have a large size set of pairwise co-prime integer matrices as matrix moduli. Also, co-prime integer matrices have applications in multidimensional sparse sensing [9]. Unfortunately, due to the non-commutativity of matrices, it is not obvious to construct a family of pairwise co-prime integer matrices of any dimension, although co-primality of integer matrices has been studied in [8] mostly for dimensions $2$ and $3$. Interestingly, a family of pairwise co-prime integer matrices was most recently obtained in [17] for any dimension.
It is known that any set of prime integers are pairwise co-prime. This holds for any prime algebraic integers as well. Therefore, it is usually an easy way to construct families of pairwise co-prime elements by simply constructing families of prime elements, such as prime integers. This motivates us to use prime integer matrices to characterize and construct pairwise co-prime integer matrices that are also prime.
In this paper, we first recall prime integer matrices and some of their basic properties presented in [2]. We then investigate co-prime and pairwise co-prime integer matrices. We show that, similar to prime integers, two different prime integer matrices are co-prime. We then present the Hermite normal forms [14] for prime integer matrices, and show that two prime integer matrices with the same determinant absolute value but different Hermite normal forms are co-prime. Thus, it is easy to construct families of pairwise co-prime integer matrices by simply contructing families of prime integer matrices with different determinant absolute values that are prime and/or with the same determinant absolute value that is prime but different Hermite normal forms. This characterizes all pairwise co-prime integer matrices that are all prime. It is interesting to note that none of the pairwise co-prime integer matrices in the construction in [17] is prime when the dimension is above $1$. Also note that a different kind of primality for matrices was studied in [12, 13] for a given set of matrices.
Since Gaussian integers are special $2\times 2$ integer matrices, interestingly, Gaussian primes and prime $2\times 2$ integer matrices are not exactly the same. When both real and imaginary parts of Gaussian integers are not 00, Gaussian primes and prime $2\times 2$ integer matrices are the same. Otherwise, Gaussian primes may not be prime $2\times 2$ integer matrices. However, two Gaussian integers are co-prime if and only if their equivalent $2\times 2$ integer matrices are co-prime. It is well-known that any Gaussian integer has a unique Gaussian prime factorization [3]. In contrast, an integer matrix can be unqiuely factorized to a product of prime integer matrices when the order of the prime integer matrices in the product is not considered. Note that the order of primes in a product is not a problem if the primes are communtative, such as Gaussian primes (or their equivalent $2\times 2$ integer matrices).
This paper is organized as follows. In Section II, we study prime integer matrices and their properties. We show their Hermite normal forms and characterize all pairwise co-prime integer matrices that are all prime. In Section III, we present some connections between Gaussian primes and prime $2\times 2$ integer matrices. In Section IV, we conclude this paper.
II Definitions and Properties of Prime Integer Matrcies
This paper considers integer matrices of size $D\times D$ for a positive integer $D>1$, where all elements of matrices are integers. Some notations are as follows. $\mathbb{Z}$ denotes the set of all integers. All vectors and matrices are $D$ dimensional integer vectors and $D\times D$ dimensional integer matrices, respectively, unless otherwise specified. $I$ is the $D\times D$ identity matrix, and 00 is also the all 00 matrix or vector. $\det(A)$ denotes the determinant of matrix $A$. And diag stands for a diagonal matrix.
Below we introduce some necessary concepts on integer matrices and for details, see, for example, [5, 6].
Unimodular matrix: A square matrix is called unimodular if its determinant is $1$ or $-1$. Otherwise, it is called non-unimodular.
Divisor and greatest common left divisor (gcld): A nonsingular integer matrix $A$ is a left divisor of a matrix $M$ if $A^{-1}M$ is an integer matrix. If $A$ is a left divisor of each of all $L\geq 2$ matrices $M_{1},M_{2},\dots,M_{L}$, it is called a common left divisor (cld) of $M_{1},M_{2},\dots,M_{L}$. Moreover, if any other cld is a left divisor of $A$, then $A$ is a greatest common left divisor (gcld) of $M_{1},M_{2},\dots,M_{L}$, denoted by gcld$(M_{1},M_{2},\dots,M_{L})$.
[…]
Multiple and least common right multiple (lcrm): A nonsingular matrix $A$ is a right multiple of a matrix $M$, if there exists a nonsingular matrix $P$ such that $A=MP$. If $A$ is a right multiple of each of all $L\geq 2$ matrices $M_{1},M_{2},\dots,M_{L}$, $A$ is called a common right multiple (crm) of $M_{1},M_{2},\dots,M_{L}$. Matrix $A$ is a least common right multiple (lcrm) of $M_{1},M_{2},\dots,M_{L}$, if any other crm of them is a right multiple of $A$, denoted by lcrm$(M_{1},M_{2},\dots,M_{L})$. From this definition, one can see that if $M$ is an lcrm of matrices $M_{1},M_{2},\dots,M_{L}$, then $MU$ is also an lcrm of $M_{1},M_{2},\dots,M_{L}$ for any unimodular matrix $U$. It is not hard to see that if two matrices $A$ and $B$ are both lcrm of the same set of matrices $M_{1},M_{2},\dots,M_{L}$, then $A=BU$ for some unimodular matrix $U$. In addition, for any groups of $D\times D$ integer matrices $M_{1,1},\cdots,M_{1,L_{1}},\cdots,M_{k,1},\cdots,M_{k,L_{k}}$, we have
$\text{lcrm}(M_{1,1},\cdots,M_{1,L_{1}},\cdots,M_{k,1},\cdots,M_{k,L_{k}})$ (1)
$=$ $\text{lcrm}(\text{lcrm}(M_{1,1},\cdots,M_{1,L_{1}}),\cdots,$
$\quad\quad\quad\quad\text{lcrm}(M_{k,1},\cdots,M_{k,L_{k}})).$
We next introduce prime matrices [2], Chapter 14.
Definition 1
[2] A nonsingular matrix $A$ is called left prime (or simply prime) if it cannot be factorized to a product of two non-unimodular matrices. If $A$ is prime, then $A$ is called a prime matrix.
For the consistency, $1$ is considered as a conventional prime integer, which corresponds to a unimodular matrix treated as a prime matrix in this paper. It is clear that if $A$ is prime, then $AU$ and $UA$ are both prime for any unimodular matrix $U$. Since the non-commutativity of matrices and we only consider one side primality, i.e., the left primality. For this reason, in this paper, a matrix $A$ and its right multiples $AU$ for unimodular matrices $U$ are the same (or in an equivalent class). In other words, matrix $A$ and matrix $AU$ for any unimodular matrix $U$ are indistinguishable. Clearly, matrices with different determinant absolute values are different. In this sense, all the lcrm of a set of matrices $M_{1},M_{2},\dots,M_{L}$ are the same, i.e., the lcrm of a given set of matrices is unique.
Proposition 1
Two different prime matrices are co-prime.
Proof. Let $A$ and $B$ be two different prime matrices. If they are not co-prime, then their gcld matrix $C$ is not a unimodular matrix and $A=CA_{1}$ and $B=CB_{1}$ for two integer matrices $A_{1}$ and $B_{1}$. Since $A$ and $B$ are different, either $A_{1}$ or $B_{1}$ is not unimodular, which contradicts with the assumption that both $A$ and $B$ are prime. q.e.d.
Proposition 2
[2] A matrix is prime if and only if its determinant absolute value is a conventional prime integer.
Its proof is not hard and can be found in [2].
With Prop. 2, we immediately have the following corollary.
Corollary 1
If $A$ is a prime matrix, then its Smith form is $A=U\mbox{diag}(1,1,\cdots,1,\lambda_{D})V$ where $U$ and $V$ are two unimodular matrices and $\lambda_{D}$ is a conventional prime integer.
From Prop. 1, we know that a set of different prime matrices are pairwise co-prime. Then, from Prop. 2, we have the following corollary.
Corollary 2
Let $M_{1},...,M_{L}$ be $L$ different matrices, i.e., no two of them only differ by a unimodular matrix factor. If their determinant absolute values are conventional prime numbers, then they are pairwise co-prime.
For MD-CRT [15, 16] that have applications in multidimensional signal processing, it is important to have various families of pairwise co-prime integer matrices as mentioned earlier. From Corollary 2, one can easily construct families of pairwise co-prime integer matrices by contructing matrices with different determinant absolute values that are all prime integers. With this in mind, one might want to ask what happens to the matrices with the same determinant absolute value that is a prime. In this case, any of them is prime but can they be pairwise co-prime?
We next provide a complete answer to this question by using Hermite normal form of integer matrices (Chapter 14 of [2] and Chapter 14 of [14]) where is says that any matrix has a unique Hermite normal form. Combined with the primality of a matrix with the result in Corollary 1, we have the following proposition.
Proposition 3
For a prime matrix $A$ that is not unimodular, its unique Hermite normal form $H=(h_{mn})$ in $A=HU$ for some unimodular matrix $U$ satisfies:
  • 1)
    $H$ is a lower triangular matrix, i.e., $h_{mn}=0$ if $n>m$;
  • 2)
    There is one and only one diagonal element that is not $1$, i.e., there exists one and only one $m_{0}$, $1\leq m_{0}\leq D$, such that $h_{m_{0}m_{0}}=\lambda_{D}>1$ and $\lambda_{D}$ is a prime integer, and all $h_{mm}=1$ if $m\neq m_{0}$;
  • 3)
    For this $m_{0}$ and any $m$th row of $H$ with $m\neq m_{0}$, all the elements of the $m$th row left to the one $h_{mm}=1$ on the diagonal are all 00, i.e., $h_{mn}=0$ for any $1\leq n<m$;
  • 4)
    All the elements $h_{m_{0}n}$ of the $m_{0}$th row left to the diagonal $h_{m_{0}m_{0}}$ are nonnegative and strictly less than $h_{m_{0}m_{0}}=\lambda_{D}$, i.e., $0\leq h_{m_{0}n}<h_{m_{0}m_{0}}$ for $1\leq n<m_{0}$.
The above Hermite normal form $H$ is shown as
$H=\left(\begin{array}[]{ccccccc}1&\cdots&0&0&0&\cdots&0\\ &\ddots&&\vdots&\vdots&&\vdots\\ 0&\cdots&1&0&0&\cdots&0\\ a_{1}&\cdots&a_{k}&\lambda_{D}&0&\cdots&0\\ 0&\cdots&0&0&1&\cdots&0\\ \vdots&&\vdots&\vdots&&\ddots&\\ 0&\cdots&0&0&0&\cdots&1\end{array}\right),$ (2)
where $k\geq 0$ and $0\leq a_{1},...,a_{k}<\lambda_{D}$ are all integers and do not appear if $m_{0}=1$.
From the above Hermite normal form $H$ in (2), one can see that there is only one row, i.e., the $m_{0}$th row $h_{m_{0}n}$, in $H$ that has one or more elements greater than $1$, and all the other rows have only one $1$ in each row and all the other elements are 00. One example of such a Hermite normal form $H$ for $D=4$ is
$\left(\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\\ 2&0&3&0\\ 0&0&0&1\end{array}\right).$
From the above Hermite normal form of a prime matrix, one difference with the Smith form in Corollary 1 is that the element $h_{m_{0}m_{0}}=\lambda_{D}>1$ in the Hermite normal form does not have to be located in the last row of the matrix as in the Smith form in Corollary 1. Since the Hermite normal form is unique for a matrix, we have the following pairwise co-primality of the prime matrices with the same determinant absolute values.
Proposition 4
All prime matrices that have the same determinant absolute value but different Hermite normal forms (2) are different and pairwise co-prime.
Proof. Let $M_{1}$ and $M_{2}$ be two prime matrices and $|\det(M_{1})|=|\det(M_{2})|$. If $M_{1}$ and $M_{2}$ are not co-prime, then there exists a unimodular matrix $U$ such that $M_{1}=M_{2}U$. This implies that $M_{1}$ and $M_{2}$ have the same Hermite normal form and thus contradicts with the assumption. q.e.d.
As we can see from the Hermite normal form $H$ in (2), for any prime integer $\lambda_{D}>1$, its different positions $m_{0}$ on the diagonal and different values of $a_{1},...,a_{k}$ all produce different Hermite normal forms and therefore different prime matrices, i.e., they are all pairwise co-prime, from Props. 1 and 4. Thus, from Corollary 2 and Prop. 4, one is able to construct many families of pairwise co-prime matrices with different determinant absolute values that are prime and with different Hermite normal forms but the same determinant absolute value that is a prime. The above results have also characterized all pairwise co-prime integer matrices that are all prime. Interestingly, a family of pairwise co-prime integer matrices of any dimension, none of which is prime when the dimension is more than $1$, was obtained in [17].
For the MD-CRT, similar to the conventional CRT, the range of the uniquely determineable integer vectors from their vector remainders depends on an lcrm of the matrix moduli. Similar to the conventional integer case, we have the following proposition.
Proposition 5
If $L$ different matrices $M_{1},M_{2},...,M_{L}$ are prime and commutative, i.e., $M_{l}M_{m}=M_{m}M_{l}$, $1\leq l,m\leq L$, then their product is their lcrm, i.e., $M_{1}M_{2}\cdots M_{L}=\mbox{lcrm}(M_{1},M_{2},...,M_{L})$.
Proof. From Prop. 1, we know that matrices $M_{1},M_{2},...,M_{L}$ are pairwise co-prime. For any two co-prime matrices $M_{l}$ and $M_{m}$ with $l\neq m$, since they are commutative, from [7], i.e., the proposition 3 in [15], we know that $M_{l}M_{m}$ is their lcrm, i.e., lcrm$(M_{l},M_{m})=M_{l}M_{m}$. Then Prop. 5 is proved by the lemma 2 in [15] and (1). q.e.d.
Although it is not easy to study commutative matrices of a high dimension in general, it is not hard to see that for $2\times 2$ matrices, commutative matrices can be always represented by
$A(a,b)=\left(\begin{array}[]{cc}a&-\alpha b\\ b&a-\beta b\end{array}\right)$ (3)
in case $b\neq 0$, where $\alpha$ and $\beta$ are two fixed constants. It is easy to check that two matrices $A(a_{i},b_{i})$, $i=1,2$, in (3) are commutative. Thus, for any set $\mathbb{S}$ of matrices $A(a,b)$ with either different determinant absolute values that are prime or the same determinant absolute value that is prime but different Hermite normal forms are both pairwise co-prime and commutative. Therefore, by Prop. 5 their lcrm is their product that determines the range of the uniquely reconstructable integer vectors of dimension $2$ from their vector remainders modulo the $2\times 2$ matrices in $\mathbb{S}$ [15].
Interestingly, algebraic integers in a quadraic algebraic number field with a minimal polynomial $r_{2}x^{2}+r_{1}x+r_{0}$ can be equivalently represented by $2\times 2$ matrices of the form in (3) with $\alpha=r_{0}/r_{2}$ and $\beta=r_{1}/r_{2}$, where $b$ can be 00, [10, 11]. Since algebraic integers are all complex numbers and therefore naturally commutative. When $r_{2}=1,r_{1}=0,r_{0}=1$, the above algebraic integers become Gaussian integers that will be especially studied in next section.
III Connection with Gaussian Integers and Gaussian Primes
In this section, we investigate and provide some connections between $2\times 2$ integer matrices with Gaussian integers.
Gaussian integers are represented by $a+jb$ where $a$ and $b$ are conventional integers in $\mathbb{Z}$ and $j=\sqrt{-1}$. The ring of Gaussian integers is denoted by $\mathbb{Z}[j]$. A Gaussian integer $z=a+jb$ can be equivalently represented by the following $2\times 2$ integer matrix
$z=\left(\begin{array}[]{cc}a&-b\\ b&a\end{array}\right),$ (4)
which is called the matrix representation of Gaussian integer $z$.
A Gaussian integer is a unit if its absolute value is $1$. It is then not hard to see that a Gaussian integer is a unit if and only if its matrix representation (or simply matrix) in (4) is a unimodular matrix. A Gaussian integer is prime if it cannot be factorized to a product two non-unit Gaussian integers. When a Gaussian integer is prime, we call it a Gaussian prime. In the following, interestingly we will see that this primality of Gaussian integers is different from that of $2\times 2$ integer matrices defined in Section II in general.
First it is well-known [3] that, a Gaussian integer $a+jb$ with $ab\neq 0$ is prime if and only if its norm, i.e., $a^{2}+b^{2}$, is a conventional prime integer. Since $a^{2}+b^{2}$ is also the determinant absolute value of its matrix representation in (4), from Prop. 2 this primality is the same as that of integer matrices in Section II. However, a Gaussian integer $z=a+jb$ with $ab=0$ is prime if and only if $|a|$ or $|b|$ is a conventional prime integer that is equal to $3$ modulo $4$. On the other hand, for a Gaussian integer $z=a+jb$ with $ab=0$, its matrix representation is either
$\mbox{diag}(a,a)=\left(\begin{array}[]{cc}a&0\\ 0&a\end{array}\right)\mbox{ or }\left(\begin{array}[]{cc}0&-b\\ b&0\end{array}\right)$
that is not prime if $z$ is not a unit. For example, $3$ is a Gaussian prime, while its matrix representation $\mbox{diag}(3,3)$ is not prime. From the above analysis, we have the following proposition.
Proposition 6
For a Gaussian integer $z$, if its matrix representation is prime, then it is a Gaussian prime, and the opposite may not be true.
In other words, matrix primality implies Gaussian primality, and Gaussian primality may not imply matrix primality. One reason might be as follows. Since in terms of the matrix primality, a matrix can be equivalently converted to a diagonal matrix of the form in Corollary 1 by using the Smith form decomposition that may not always preserve the matrix representation structure (4) for a Gaussian integer in the decompsition steps. This means that the Smith form in Corollary 1 for a prime matrix may not apply to a Gaussian prime, i.e., a Gaussian prime does not have to have the Smith form in Corollary 1.
We next want to see the co-primality of two matrices and two Gaussian integers. First two Gaussian integers are called co-prime if they do not have any non-unit common factor. Although the primalities of a Gaussian integer and its matrix representation are not equivalent, we will show below that the co-primalities of two Gaussian integers and their matrix representations are equivalent.
Proposition 7
Two Gaussian integers are co-prime if and only if their matrix representations are co-prime in the integer matrix sense.
Proof. We first note that for both integer matrices and Gaussian integers, Bezout’s identity holds. Let $z_{i}$ have matrix representations $Z_{i}$ in (4) for $i=1,2$. From Bezout’s identity, $z_{1}$ and $z_{2}$ are co-prime if and only if there exist Gaussian integers $y_{1}$ and $y_{2}$ such that $z_{1}y_{1}+z_{2}y_{2}=q$ for a unit $q$. Let $Y_{1},Y_{2},Q$ be the matrix representations of $y_{1},y_{2},q$ in (4), respectively. Then, we have $Z_{1}Y_{1}+Z_{2}Y_{2}=Q$ and vice versa. Since $q$ is a unit if and only if $Q$ is unimodular, $z_{1}y_{1}+z_{2}y_{2}=q$ for a unit $q$ if and only if $Z_{1}Y_{1}+Z_{2}Y_{2}=Q$ for a unimodular matrix $Q$ if and only if $Z_{1}$ and $Z_{2}$ are co-prime by Bezout’s indentity for integer matrices [9]. q.e.d.
It is well-known [3] that any Gaussian integer has a unique prime factorization, i.e., any Gaussian integer can be uniquely factorized to a product of some prime factors. This is also true for matrix prime factorization if the order of the prime factors in the product is not considered.
Proposition 8
A non-singular matrix $A$ can be uniquely factorized to a product of prime matrices called prime factors, if the order the prime factors in the product is not considered.
Proof. Let $A=IP_{1}P_{2}\cdots P_{L}$ be a prime factorization of matrix $A$, where all $P_{i}$ are prime. For each $P_{i}$, let $P_{i}=U_{i}\Lambda_{i}V_{i}$ be its Smith form decomposition with unimodular matrices $U_{i},V_{i}$, where $\Lambda_{i}$ is a diagonal integer matrix, $i=1,2,...,L$. Since all $P_{i}$ are prime, from Corollary 1, $\Lambda_{i}=\mbox{diag}(1,1,...,1,\lambda_{i})$ for some conventional prime integers $\lambda_{i}$. Since $|\det(A)|=\prod_{i}^{L}\lambda_{i}$ whose factorization is unique, and $IW_{1},\Lambda_{1}W_{2},...,\Lambda_{L}W_{L+1}$ are the same as $I,\Lambda_{1},...,\Lambda_{L}$ for all unimodular matrices $W_{1},W_{2},...,W_{L+1}$, respectively, the above factorization of matrix $A$ is unique if the order of the factors in the product is not considered. q.e.d.
A similar integer matrix factorization can be found in [2]. Although the above connection is only studied for Gaussian integers, similar studies may apply to other algebraic integers in other algebraic number fields.
IV Conclusion
In this paper, we have applied prime integer matrices to construct and characterize all pairwise co-prime integer matrices that are all prime. For instance, two prime integer matrices with different determinant absolute values that are prime are co-prime, and two prime integer matrices with the same determinant absolute value that is prime but different Hermite normal forms are co-prime. These pairwise co-prime integer matrices may have applications in multidimensional CRT and multidimensional sparse sensing. We have also investigated the connections between prime integer matrices and Gaussian primes.